# FB-RAG: Improving RAG with Forward and Backward Lookup

### v2

@@ Appendix B Datasets
Our experiments are based on datasets from two popular benchmarks consisting long context lengths - LongBench bai-etal-2024-longbench and Bench zhang-etal-2024-bench. QA tasks (NarrativeQA, Qasper, MultifieldQA, HotpotQA, 2WikiMultihopQA, MuSiQue, and EN.QA) take a query and a context as input, with the goal of generating a concise answer. The summarization task (QMSum) requires generating a free-form summary based on the given query and context. For the MCQ task (EN.MC), the input additionally includes a set of choices, and the task is to choose the correct choice that answers the input query based on the provided context. We present key statistics for these datasets in Table 3.
@@
Bench zhang-etal-2024-bench | ||
@@ Appendix C Experiment Design
We provide additional experimental design details in this section to promote reproducibility. We further plan to release our code on acceptance.

### v3

@@ 7 Conclusion
Our findings also provide clear guidance on when to use FB-RAG. For applications with similar setup as ours, Ours-F (forward-lookup only) should be preferred over Ours-FB (forward and backward) given its superior performance in our experiments. The choice between Ours-F and OP-RAG depends on the specific performance versus latency requirements:
Fixed performance target: As shown in Figure 2 (bottom), OP-RAG may be suitable for lower performance targets, but for higher targets, Ours-F achieves them with lower latency.
Fixed latency budget: For low latency budgets, OP-RAG is preferable, but as the budget increases, Ours-F delivers superior performance.
This highlights a key advantage of FB-RAG for long-text applications. When the input context is large (e.g., 16k or 24k), standard RAG with a large generator (e.g., 70B model) incurs high latency. In such cases, FB-RAG’s strategy of using a light-weight model for forward-lookup allows it to achieve higher performance while simultaneously reducing overall latency.
@@ Appendix B Datasets
Our experiments are based on datasets from two popular benchmarks consisting long context lengths - LongBench Bai et al. (2024) and Bench Zhang et al. (2024). QA tasks (NarrativeQA, Qasper, MultifieldQA, HotpotQA, 2WikiMultihopQA, MuSiQue, and EN.QA) take a query and a context as input, with the goal of generating a concise answer. The summarization task (QMSum) requires generating a free-form summary based on the given query and context. For the MCQ task (EN.MC), the input additionally includes a set of choices, and the task is to choose the correct choice that answers the input query based on the provided context. We present key statistics for these datasets in Table 3.
@@
| Bench Zhang et al. (2024) | ||
@@ Appendix C Experiment Design
We provide additional experimental design details in this section to promote reproducibility.
@@ D.4 Varying the number of samples used in Stage II of FB-RAG
D.5 In-Context Learning experiments
Section 5 discusses additional analysis where we combine the techniques developed in this work with the complementary methods proposed by Wei et al. (2024). We present the results from these experiments in Table 9.
@@ D.5 In-Context Learning experiments
| Method | Avg | Narr | Qasp | Mult | Hotp | 2Wiki | Musi | QMSum |
| Results copied from Table 1 in the main paper | ||||||||
| Vanilla | 44.19 | 25.01 | 49.31 | 53.41 | 60.91 | 58.84 | 37.32 | 24.51 |
| Augmented to Vanilla RAG | ||||||||
| Reason-then-Answer | 33.42 | 19.01 | 46.08 | 45.34 | 43.78 | 29.6 | 26.49 | 23.61 |
| Few-Shot Demo. w/ Instruction | 41.59 | 21.24 | 46.49 | 50.54 | 54.49 | 56.24 | 37.75 | 24.36 |
| INSTRUCTRAG-ICL | 42.95 | 23.06 | 47.55 | 53.74 | 54.8 | 61.47 | 35.93 | 24.08 |
| Augmented to Ours-F (6k 1.5k) | ||||||||
| Reason-then-Answer | 37.71 | 27.74 | 48.48 | 45.33 | 46.61 | 37.06 | 35 | 23.75 |
| Few-Shot Demo. w/ Instruction | 46.76 | 27.26 | 49.73 | 50.52 | 58.21 | 64.52 | 52.67 | 24.42 |
| INSTRUCTRAG-ICL | 48.79 | 28.2 | 48.99 | 55.52 | 63.7 | 69.42 | 51.51 | 24.22 |